### v2

# DECO: Life-Cycle Management of Production-Scale Copilots [Industry]

Abstract.
Software engineers frequently grapple with the challenge of accessing disparate documentation and telemetry data, including TroubleShooting Guides (TSGs), incident reports, code repositories, and various internal tools developed by multiple stakeholders. To enhance the efficiency of on-call engineers (OCEs) and streamline their daily workflows, we introduced DECO—a comprehensive framework for developing, deploying, and managing copilots tailored to improve productivity in engineering routines. This paper details the design and implementation of the DECO framework, emphasizing its innovative NL2SearchQuery functionality and a lightweight agentic framework. These features support efficient and customized retrieval-augmented-generation (RAG) algorithms that not only extract relevant information from diverse sources but also select the most pertinent skills in response to user queries. This enables the addressing of complex technical questions and provides seamless, automated access to internal resources. Additionally, DECO incorporates a robust mechanism for converting unstructured incident logs into user-friendly, structured guides, effectively bridging the documentation gap. Since its launch in September 2023, DECO has demonstrated its effectiveness through widespread adoption, enabling tens of thousands of interactions and engaging hundreds of monthly active users (MAU).
*Work done while interning at Microsoft.
PVLDB Reference Format:
Yiwen Zhu, Mathieu Demarne, Kai Deng, Wenjing Wang, Nutan Sahoo, Divya Vermareddy, Hannah Lerner, Yunlei Lu, Swati Bararia, Anjali Bhavan, William Zhang, Xia Li, Katherine Lin, Miso Cilimdzic, Subru Krishnan. PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
1. Introduction
The daily activities of software developers encompass a broad spectrum of tasks that extend beyond mere development activities such as coding, debugging, and testing (Meyer et al., 2019). Predominantly, engineers allocate the majority of their efforts to ancillary tasks including code review, documentation, and on-call responsibilities (Gonçalves et al., 2011). These duties necessitate extensive manual searches across disparate resources—ranging from product documentation and troubleshooting guides to telemetry data, system logs, code repositories, and team reports. Moreover, engineers invest considerable effort into navigating internal tools that are essential for their daily operations during the arduous task of managing such varied information sources. The challenge is compounded when critical institutional knowledge is isolated or hoarded by senior engineers, resulting in significant knowledge gaps when these key individuals are absent or depart from the organization.
Recent advancements in large language models (LLMs), such as GPT (Achiam et al., 2023), LLaMA (Touvron et al., 2023), and Gemini (DeepMind, 2024), have emerged as powerful tools in this context. These models offer significant potential to enhance the engineers’ efficiency by automating mundane tasks, providing swift access to relevant information, and bridging knowledge gaps. The deployment of such AI tools not only facilitates more efficient workflow but also ensures centralized access to a diverse array of resources, thereby streamlining the decision-making process for software development teams.
Introduction to DECO—a framework for developing, deploying, and managing copilots
In this paper, we introduce DECO, a framework for developing, deploying, and managing copilot applications designed to support the continuous evolution of a software development team’s knowledge base while simplifying access to various internal tools. Copilots configured with DECO streamline daily tasks, improve incident management, and provide engineers with a unified access point to a wide range of internal resources and tools. DECO focuses on four key areas: (1) a Generalized Development Platform, enabling contributors to easily create and deploy new copilots; (2) Skills/Agent Integration, allowing teams to add new capabilities that incorporate additional knowledge sources or expose skills for accessing internal tools through a chat interface; (3) Fine-Tuned Retrieval Algorithms, designed for efficiency and accuracy in production environments; and (4) Streamlined End-to-End Deployment and Lifecycle Management, offering a self-hosting model with fully automated deployment pipelines, continuous monitoring, and testing capabilities to support debugging and ongoing improvements. Additionally, DECO supports an extensible development platform that enables integration with internal tools—such as debugging utilities and monitoring systems—frequently used in the daily tasks of engineers.
Figure 1. Monitoring dashboard for Azure Data services
Using DECO, dozens of teams across various divisions at Microsoft have successfully deployed copilots tailored to their specific roles. Within the Azure Data organization—which includes services such as Azure SQL DB, Microsoft Fabric, and Microsoft Synapse—we have recorded tens of thousands of messages and over 500 monthly active users (MAU) [^1] Many users are on-call engineers, typically working in rotation. (see Figure 1). Across approximately 1,500 monthly sessions, conversations average 2.2 rounds, indicating typical multi-turn interactions. From the several hundred feedback entries received (rated in stars), the average rating is 3.6. Interviews with on-call engineers (OCEs) reveal that DECO not only supports routine tasks but also significantly reduces the time required for incident triaging—by approximately 10–20 minutes per incident.
Introduction to DECO
DECO introduces the following innovations:
Generalized Framework. Traditional development of copilot applications typically involves iterative adjustments and prompt tuning through trial and error to attain satisfactory performance. In contrast, DECO introduces an “autonomous mode” that facilitates self-onboarding for tenants, adopting a “one-size-fits-all” approach posing a higher level of development sophistication.
Knowledge Gap. Recognizing the challenges of maintaining high-quality documentation and its critical role in providing accurate context to large language models (LLMs), DECO implements a robust mechanism to transform large volumes of incident logs into user-friendly, structured guides, enriching the knowledge base.
Response Quality. To enhance the quality of the copilot’s responses, we implemented several key improvements:
DECO employs an agentic framework for skill and agent selection, integrating advanced retrieval algorithms to optimize both accuracy and functionality. The most common issues with the copilot are not due to hallucinations (see Section 8.2, where only 0.1% of responses are identified as having grounding issues), but rather document relevancy problems. This challenge becomes particularly pronounced when managing a large inventory of skills and documentation segments. Furthermore, the documentation sources for DECO copilots often consist of highly technical materials with similar content, which further complicates the retrieval process.
To manage diverse functionalities and enhance response accuracy, DECO employs an agentic framework to select the appropriate agent and corresponding skills for addressing user queries. Compared to using a single planner to choose from a larger set of skills, this sequential selection approach significantly improves skill selection accuracy. Additionally, we develop an an innovative “NL2SearchQuery” method to enhance document retrieval from documentation sites, Incident Management (IcM) databases, and other text-based sources. For code search, we introduce the Language-Agnostic RA (LA-RAG), which preprocesses large code repositories containing multiple languages while preserving code component relationships. The solutions are not only lightweight (requiring no fine-tuning or post-training) but are also supported by existing commercial indexing services, such as Azure AI Search (Microsoft, 2024b).
DECO incorporates a comprehensive evaluation system alongside telemetry data collection pipelines to facilitate online monitoring and debugging. Following deployment across multiple tenants, we continuously monitor response quality through a robust evaluation framework. This system assesses copilot performance using diverse metrics, including answer relevance as judged by LLMs and user feedback. Additionally, specialized evaluators are implemented to verify the accuracy of individual modules within the copilot workflow and identify opportunities for improvement. Hosted on Azure, DECO features a comprehensive telemetry infrastructure that not only tracks copilot response quality but also supports efficient logging for rapid issue triaging.
Contribution
The main contributions of DECO are as follows:
  • An end-to-end pipeline for the development, deployment, and management of production-grade copilots that support software engineers’ daily routines and incident resolution.
  • Conversion of incident databases into curated reports, addressing gaps in existing documentation.
  • An efficient retrieval algorithm utilizing an NL2SearchQuery module and a Language-Agnostic RAG (LA-RAG) for document search.
  • A lightweight agentic framework for skill/agent selection, enabling access to documentation, internal toolkits, and code, thereby enhancing chat quality.
  • A robust evaluation and monitoring system to drive continuous improvement and streamline error triaging.
The remainder of the paper is organized as follows: Section 2 covers background information and data sources. Section 3 outlines the overall system design. Section 4 describes the data preprocessing pipelines for knowledge base preparation. Section 5 details the backend infrastructure, and Section 6 explains the advanced document retrieval algorithms. Section 7 presents the frontend components. Section 8 describes the evaluation framework and experimental results. Section 9 discusses related work, and Section 10 concludes the paper.
2. Background
The primary objective of DECO is to provide a framework for setting up copilots that offer engineers a time-efficient, interactive experience for managing daily tasks and on-call responsibilities. DECO effectively bridges gaps between product telemetry, documentation, on-call duties, code development resources, and team-specific toolkits.
2.1. Data Sources
By default, DECO incorporates resources from (1) internal documentation repositories, such as troubleshooting guides; (2) historical incident databases; (3) code repositories; and (4) internal Stack Overflow. Teams can also add additional skills to integrate more sources as needed.
Documentation Site
Microsoft maintains an internal documentation site that serves as a central hub for critical information, including engineering documentation and troubleshooting guides relevant to various products and services.
IcM
Incident data, including information such as title, summary, source, and owning team, is stored in a dedicated database (Chen et al., 2020). This database also captures all communications between engineers, including investigation notes. These records serve not only as a comprehensive documentation resource for incident triaging but also provide invaluable guidance for broader engineering practices.
Code Base
For questions related to code, we process the original code repository to create indexes that facilitate easy retrieval of relevant functions, classes, statements, and other code elements.
Stack Overflow
Similar to IcM databases, Stack Overflow serves as an invaluable resource for Microsoft engineers, offering a wealth of community-driven knowledge and problem-solving techniques. Discussions from these forums are extracted and summarized into reports for future reference. This section provides a comprehensive overview of DECO’s architecture and design decisions.
Figure 2. DECO architecture
3. Architecture
DECO comprises 4 primary components: Data Preprocessing, Backend Services, Frontend Services, and Evaluation (see Figure 2):
Data Preprocessing: The data preprocessing pipelines are orchestrated in the offline process, executed at user-defined intervals to account for information updates frequently. They draw inputs from various data sources, such as Git, Wikis, and Kusto databases, to generate processed documents stored in the cloud, such as Azure Blob Storage, for online retrieval. Multiple document search indexing services (such as Azure AI Search (Microsoft, 2024b)) are deployed to index the processed documents, enabling seamless access through API calls from downstream backend (online) modules.
Backend Services: The backend orchestrates the agentic framework using PromptFlow (Microsoft, 2024k) for the online interaction with users. It enables straightforward deployment, API and flow monitoring, and greatly reduces maintenance complexity for developers. The workflow can be easily deployed in the cloud, serving as a stateless REST API through managed online endpoint (Microsoft, 2024d) in Azure Machine Learning (Microsoft, 2024c), optimized for real-time, high-volume, low-latency inference requests, and equipped with extensive logging and debugging capabilities, and can be easily scaled to support hundreds of users. Sections 5 and 6 provide further details on these modules.
Frontend Services: The frontend manages user authentication and authorization and handles user interactions. It manages session memory, building the full chat history and sending it to the backend’s (stateless) REST API to enable a continuous conversation experience. Session state and extracted memory data are stored in the cloud, while Azure Log Analytics automatically stores telemetry data for easy monitoring and review from the API calls. This data is later used for feedback learning (Anonymous, 2024b) and evaluation (see Section 8). A web application is deployed with two main user interfaces: a web-based platform and a Microsoft Teams integration, including group chat and session-sharing functionalities.
Compared to many existing agentic frameworks where session and memory management are handled within the same backend (e.g., AutoGen (Wu et al., 2023) or Letta (Packer et al., 2023)), implementing a stateless REST API in the backend to operate independently of session state management and authentication from the frontend, offering two key advantages in production:
  • Increased development agility through simplified implementation due to decoupling, and
  • Seamless backend updates for the copilot, with the flexibility to use multiple backend endpoints to manage traffic, as session state is handled by the frontend.
Evaluation: A comprehensive assessment of the entire system is conducted to monitor and evaluate real user interactions as well as individual modules, ensuring a high-quality chat experience. The evaluation pipeline supports: (1) Pull Request checks and daily monitoring to track response quality and prevent regressions; (2) leveraging user feedback to construct golden datasets that guide module development and improve response quality; and (3) using user feedback to enhance documentation retrieval, such as through adaptive re-ranking (Anonymous, 2024b).
4. Data Preprocessing
In this section, we describe the offline preprocessing pipeline for three data sources: IcM (historical mitigated incidents), documentation, and code repositories. These pipelines build the knowledge base (index) to support context-retrieval skills in the backend, such as get_tsg, get_icm, and get_code. Additional pipelines can be created using the same infrastructure to integrate information from other sources, such as Stack Overflow or internal email exchanges.
4.1. Common Infrastructure
DECO provides a unified data preprocessing infrastructure that standardizes ingestion, transformation, and indexing across all pipelines. This common infrastructure includes the following modules, each of which can be customized upon instantiation:
  • Data Ingestion: We use secure authentication to access various sources, such as the IcM database, Git repositories, and documentation sites.
  • Data Transformation: We develop asynchronous scheduled pipelines that extract, and transform raw data to generate structured content in the cloud.
  • Metadata Enrichment: We compute of additional metadata fields, such as reference links and key attributes that can be used to enhance retrieval and re-ranking.
  • Embedding Generation: We use LLMs to generate embeddings across various fields in the structured content.
  • Data Indexing: We create of search indexes based on processed data, with deployment to provide a REST API for efficient querying.
4.2. IcM Processor
The IcM Processor extracts valuable insights from previously mitigated incidents, where engineers may have logged useful queries, mitigation steps, and commands. To better leverage this scattered data, we have decided to preprocess the information into a structured format, formulating additional documentation for the copilot. The output of this pipeline supports the get_icm skill.
Built on the common infrastructure described in Section 4.1, The IcM Processor consists of the following customized steps:
  • Data Injection: The Kusto Connector utilizes secure authentication via managed identity (Microsoft, 2024j) to access the IcM database. It extracts free-form records based on user-specified team name(s) and date range from multiple data tables. The set of IcMs to be included can be customized by different product teams.
  • Data Transformation: The transformation module employs GPT-4o models to generate structured summaries from free-form text data. This process leverages predefined summarization prompts and curated output templates to extract key fields such as summary, title, mitigation steps, and properties to enable specific lookups (see Section 6.2). Previous work has demonstrated that LLMs have effectively tackled this task (Pu et al., 2023).
  • Metadata Enrichment: Additional metadata is computed, such as a helpfulness score, reflecting the thoroughness of the generated documentation.
Figure 3 illustrates the input and output of IcM Processor. Typically, each copilot creates indexes for approximately 3,000 to 30,000 incidents.
Figure 3. Example of data preprocessing pipeline for IcM Processor
4.3. TSG Processor
The TSG Processor accesses Git repositories, splits long documents into smaller chunks, and generates embeddings for multiple fields, including document content, titles, and URLs. This enables multi-field hybrid search, improving accuracy and ranking. Additional metadata, such as document descriptions, referenced issue IDs, and base64-encoded images are also extracted. The pipeline supports advanced search algorithms, including Reverse-HyDE question embeddings (Gao et al., 2022) and techniques from (Anonymous, 2024a). We conducted a thorough evaluation of documentation retrieval algorithms and hyper-parameters for DECO in (Anonymous, 2024b).
Typically, each copilot indexes approximately 5,000 to 50,000 troubleshooting guide chunks for optimized retrieval.
4.4. Code Processor
Modern software codebases are inherently multi-language, often integrating diverse programming paradigms and a variety of file formats. Beyond source code, typical codebases include configuration files such as YAML, JSON, and XML, which play a critical role in defining system behavior but remain outside the main compilation process. This heterogeneity poses significant challenges for compiler-based indexing tools attempting to comprehensively analyze and query codebases.
Abstract Syntax Tree (AST)-based indexing methods, such as those leveraging Tree-Sitter (Brunsfeld et al., 2022), primarily focus on the syntax of individual programming languages while neglecting other essential components like configuration files, external dependencies, and runtime interactions. As a result, these methods often miss crucial relationships that span across language boundaries. Additionally, because compilation or interpretation usually occurs independently for each language (e.g., Python and C# components are processed separately), traditional indexing techniques struggle to capture the holistic behavior of multi-language systems. Additionally, conventional document chunking and vector-based retrieval approaches frequently struggle to preserve the intricate dependencies and relationships within a codebase. Even worse, these methods risk fragmenting single code segment into multiple chunks, disrupting the continuity of processed documentation and impairing the model’s ability to accurately answer complex user queries.
To address these limitations, DECO introduces the Language-Agnostic RAG (LA-RAG), which leverages LLMs’ multi-language capabilities without relying on compilation. Unlike in-memory Copilot solutions that process only open files within an Integrated Development Environment (IDE), our offline indexing mechanism enables comprehensive search across large-scale codebases containing thousands of files and millions of lines of code spanning multiple languages by structuring the codebase as a loosely coupled reference graph. For the online retrieval, our system is able to accurately determine code construct usage, ensuring context-aware retrieval beyond immediate source files.
The preprocessing pipeline is developed using the same infrastructure as discussed in Section 4. Similarly as the TSG Processor, the Code Processor access the code based from Git repositories, and tentatively splits long code files into fix-size smaller chunks. After generating code chunks, we apply the following metadata enrichment and embedding steps:
  • LLM Rechunking: To prevent functions or classes from being split across chunks, we concatenate each chunk with its five preceding and succeeding chunks, then use an LLM to extract complete code segments.
  • Component Relationships: We use LLM to identify one-hop related components and attach their corresponding code chunks, improving contextual understanding.
  • Description Augmentation: The same LLM call generates a title, a concise description, references used within the code chunk, and summaries of referenced functions or classes to enhance searchability.
  • Embedding Generation: We generate embeddings for code chunks and associated metadata, creating a structured search index for efficient retrieval across multiple fields that is used in the online phase (see Section 6).
Typically, each copilot indexes approximately 300 to 20,000 code chunks for optimized search and retrieval.
5. Copilot Backend
The backend REST API is implemented as a stateless API, requiring all necessary information, such as the full chat history as input. This design abstracts away much of the engineering overhead related to authentication and state/memory management, which is handled by our bot frontend (discussed in Section 7). The backend orchestration is created to maximize re-usability across different copilots and allow easy customization of incorporate different skills/configuration settings from different teams.
5.1. Design Principles
The DECO’s backend orchestration is guided by the following core principles:
Multi-turn interactions: The future of generative AI platforms lies in multi-agent, multi-turn copilots (Wu et al., 2023). This pattern is emerging across various existing copilots, from OpenAI’s models iteratively correcting their own generated code during sandbox execution to multi-step data-gathering workflows for complex tasks (Zheng et al., 2025).
In DECO, all iterative interactions are managed through the frontend, ensuring full transparency and user control. Past interactions are retained in history, allowing users to provide additional inputs when necessary. Unlike platforms that orchestrate multiple agents behind a single API call (see Figure 4(a)), DECO does not require custom handling mechanisms for multi-turn interactions but only execute one agent in one invocation of the backend (see Figure 4(b)), which offers better control compared to existing free-chat approaches, such as AutoGen (Wu et al., 2023).
(a) A single API call.
(b) Multi-turn interactions.
Figure 4. Frontend-backend interactions
Simplicity and Maintainability To facilitate ease of development, debugging, and extensibility, DECO adopts a modular and interpretable workflow based on a stepwise “tool-use scheme”. Execution follows a structured model where each agent sequentially invokes a subset of skills, actively interacting with the frontend. This modular execution allows engineers to inspect intermediate states, log agent outputs, and isolate failures efficiently. Since agents are loosely coupled, individual components can be updated, replaced, or extended without affecting the overall pipeline. Additionally, each execution round generates transparent logs, making it easier to trace and debug incorrect responses.
Low-Latency Execution DECO ensures real-time responsiveness by streaming intermediate outputs from agents as they execute. Unlike traditional batched execution, where users must wait for all skills to complete, our framework delivers partial responses incrementally (see Figure 4(b)). This approach enhances usability by allowing users to refine their input based on intermediate results, ensuring a dynamic and interactive experience. Furthermore, since each step displays output progressively, users can cancel an interaction at any time without losing already provided information.
Scalability and Efficient Planning Flat planners, which evaluate all skills simultaneously, become inefficient as the number of available skills grows. As the skill set expands, a flat planner must consider an exponentially increasing number of possible combinations, leading to computational inefficiencies and suboptimal execution sequences. Prior research, such as fine-tuning-based approaches (Patil et al., 2023), has attempted to mitigate this by training LLMs to generate skill execution plans. However, fine-tuning requires extensive labeled data, making it both costly and impractical in dynamic environments where new skills are frequently introduced.
To overcome these limitations, DECO employs a hierarchical agentic planning approach, dynamically selecting and sequencing agents based on the user query. Instead of treating skills as isolated decision points, the system:
  • First determines the relevant agent (e.g., troubleshooting, database analysis, code retrieval).
  • Then, based on the selected agent, determines the appropriate skills to invoke.
This hierarchical approach reduces complexity, improves efficiency, and maintains flexibility for integrating new skills, ensuring DECO remains adaptable as capabilities evolve.
5.2. Orchestration
Figure 5. Backend orchestration
Figure 5 illustrates the backend orchestration of the copilot, addressing the above consideration. In each round of backend workflow invocation, the Manager module performs the following functions:
  • Determines whether the previous invocation sufficiently answers the user’s question (Termination Decision Maker).
  • Dynamically updates the meta-plan, which specifies the sequence of agents to execute (Planner) and identifies the current agent to be executed in this round.
The Executor runs the current agent and generates the output. This output is sent to the frontend in streaming mode, automatically triggering the next workflow invocation (without user input) to execute the next agent in the meta-plan. This process continues until the Termination Decision Maker returns the termination flag. From the user interface, users can sequentially view outputs from each agent without waiting for the entire meta-plan execution to complete.
5.3. Agents and Skills
Agent
To enhance DECO’s modularity, we introduce the concept of “agents”, enabling onboarding teams to seamlessly integrate with various skills or even existing copilots through APIs. Each agent is defined by (1) a detailed description of applicable scenarios; (2) a set of skills the agent can invoke; (3) (optional) an LLM prompt that the agent uses at the end of execution to generate a response, combining all skill outputs as input context; and (4) (optional) a hint for the next agent to be invoked. Based on the setting of (3), the agent can either invoke an additional LLM call or directly return the skill output as the agent output for this invocation round.
For example, the Kusto Agent includes the skills kql.generator and kql.executor; the Root Cause Analysis Agent uses the RCA.copilot skill; the Code Agent utilizes the get_code skill; and the QA Agent employs the skills get_tsg, get_icm, etc.
Skill
Each skill extends from a common abstract class and is defined by (1) a detailed description of its functionalities; (2) a specification of its input arguments, such as a rephrased user question, which need to be extracted from the skill input; (3) the execution code (in Python), which may involve API calls, LLM interactions, or any other execution processes; and (4) (optional) skill dependencies, which impacts the execution order, ensuring that required outputs from preceding skills are available as inputs.
For example, the get_tsg skill requires a “rephrased user intent” as an argument and retrieves relevant documentation via the documentation search API. Its output consists of a list of TSG chunks, which can later be leveraged by the agent’s final LLM call to generate a response. The kql.executor skill can execute Kusto queries on behalf of users.
5.4. Manager
The Termination Decision Maker and the Planner execute in parallel during each workflow invocation round.
The Termination Decision Maker collects the chat history (i.e., the user’s question and responses from previously executed agents) and makes a termination decision using an LLM call. The prompt for this call includes detailed termination conditions and examples.
The Planner loads all available agents and compiles their names, descriptions, and hints for next agent into an agent selection prompt. An LLM generates the sequence of agents to be invoked in the meta-plan. In each round of workflow invocation, the Planner updates the meta-plan and selects the next agent for execution. In our current implementation, most user queries are resolved with fewer than three agents, and the Planner’s decisions are highly accurate.
5.5. Executor
Once the Planner selects an agent for execution, the Executor performs the following steps to “execute” the corresponding agent: Step 1: Selects the skills to execute from the agent’s available skill set and extract arguments; Step 2: Generates a skill execution plan based on defined skill dependencies (if any); Step 3: Executes the skill execution plan and collects outputs from all invoked skills; and Step 4 (Optional): Constructs the prompt based on the agent’s definition and makes a final LLM call to generate the agent’s response.
In Step 1, we compile descriptions of all available skills for the selected agent, along with their required input arguments, into the skill selector’s prompt. Few-shot examples are included to help guide the LLM in selecting the appropriate skills based on various user inputs. The output of this LLM call includes both the list of skills to invoke and the corresponding input arguments for each skill. For the skill orchestration LLM call, we use the Semantic Kernel (Microsoft, 2024g) sequential planner, integrated with Azure OpenAI function calling (Microsoft, 2024e) to restrict the output format. In Step 1, the LLM determines only the set of skills to execute. Step 2 then applies a deterministic optimization mechanism to decide the execution order and parallelism, accounting for dependency constraints. If no constraints exist, all skills are executed in parallel. For certain agents, some skills are capable of directly generating user responses and terminate this round of invocation.
Figure 6. NL2SearchQuery
6. Retrieval Algorithms
In this work we proposed lightweight retrieval strategies and re-ranking algorithms for skills such as get_icm, get_tsg, and get_code to enhance retrieval precision in production, supported by matured commercialized search services, such as Azure AI search (Microsoft, 2024b).
6.1. NL2SearchQuery
In the preprocessing pipeline, the processed data often includes several important search fields. For example, “resolve date” for IcMs, “title” and “content” for TSGs, and “reference” or “description” for code. Using an LLM, we can infer optimal matching criteria or additional filters based on user input to improve the retrieval results. For instance, queries such as “all the code related to the class SOSScheduler” or “show me recent incidents similar to Incident XXX” can be formulated as different search queries.
Since all fields are indexed, similar to a SQL query, we can search for relevant information by matching against different fields and applying different filters/predicates to pinpoint the most relevant documents. This insight informs the design of our NL2SearchQuery algorithm, as illustrated in Figure 6. Given a user query, we leverage an LLM call to extract key arguments for retrieval and construct a complete search query which consists of selected search fields, time ranges, or any other filtering criteria predefined for each retrieval skill. Multiple search queries can be executed in parallel, and the combined result set can be re-ranked using different ranking strategies to produce the final output. The following sections provide a detailed discussion of the retrieval strategies for IcM, TSG, and code respectively.
6.2. IcM Retrieval
The get_icm skill, essential for generating curated incident reports, is frequently used to respond to user inquiries about how-to guides and to search for similar incidents, thereby offering additional mitigation strategies. We’ve noticed that user queries invoking this skill typically highlight specific incident details, such as failure mechanisms or server names, which align with the mitigation steps or extractable properties of incidents (discussed in Section 4). For this task, the following arguments are extracted to form the search query:
  • User Intent: Reformulates the user query into concise search keywords, considering not only the original question but also the full chat history and previous intermediate outputs to create precise search terms. This is used as the search predicate for the matching.
  • Search Field: Specifies the field to search within, such as “summary”, “title”, “property”, or “mitigation”. For example, if the user asks for similar incidents to a given one, “title” is often a suitable option. If the query includes specific properties like server name, “property” is used.
  • Search Method: Options include hybrid search, semantic search, simple search, or vector search (Microsoft, 2024f).
  • Time Range: Inferred from the user query and extracted as a dictionary where the key is the date type and the value is the time range in days.
  • Ticket Type: Specifies the type of incident, with possible values including “LSI” (Live Site Incident), “CRI” (Customer Reported Incident), or “ALL”.
Through testing, we found that function calling with the latest GPT-4o model, supported by few-shot examples, can accurately extract all necessary arguments based on user intent. The following examples illustrate the generation of search queries:
  1. (1)
    Question: “Show me customer-reported incidents resolved by restarting the server in the last two weeks.”
    • User Intent: Customer-reported incidents resolved by restarting the server.
    • Search Field: “mitigation”.
    • Time Range: {“resolve_date”: 14}.
    • Incident Type: CRI (Customer Reported Incident).
  2. (2)
    Question: “Are there any live site incidents created in the last two days involving issues on server testserver1?”
    • User Intent: Issues on server testserver1.
    • Search Field: “property”.
    • Time Range: {“create_date”: 2}.
    • Incident Type: LSI (Live Site Incident).
Re-ranking. The initial search query typically returns a large set of incidents, for example, 20. During the re-ranking step, we calculate scores for each incident to determine the top $K$ results, such as the top 4, for the final retrieval output.
The re-ranking score for a document $d$ is calculated as follows:
(1) $P(d)=\alpha\cdot\text{IS}+\beta\cdot\text{TS}+\gamma\cdot\text{SS}$
where IS represents the information score, which evaluates the quality of the IcM summary based on data quality metrics like token length, pre-computed during the preprocessing pipeline (see Section 4). TS, the time score, assesses the relevance of incidents based on their age, with older incidents presumed less relevant and assigned lower values. SS, the source score, verifies if the retrieved data aligns with the current context, assigning a value of 1 for matches (e.g., matching team or monitor ID in property fields) and 0 otherwise. These scores are normalized and combined to re-rank the retrieved incidents, focusing on factors like information richness, timeliness, and source relevance. This process is detailed in Algorithm 1. The accuracy of our ICM retrieval algorithm significantly outperforms the default similarity search in the ICM Portal (see Section 8.1).
1:  Input: User question $q$, IcM retrieval database $D$, number of top results $K$, number of incidents to retrieve per query $N$, re-ranking score function $P(d)$
2:  Output: Top $K$ relevant documents $D^{*}$
3:  $S\leftarrow$ Generate a set of AI Search queries based on user question $q$
4:  Initialize an empty set $D^{*}$ for retrieved documents
5:  for each $s_{i}\in S$ do
6:     $D_{i}\leftarrow$ Execute $s_{i}$ to retrieve top $N$ incidents from $D$
7:     Compute the re-ranking score $P(d)$ for all $d\in D_{i}$
8:     Add $D_{i}$ to the set $D^{*}$
9:  end for
10:  Rank all documents in $D^{*}$ based on the re-ranking score
11:  $D^{*}\leftarrow$ Select top $K$ documents from the ranked list
12:  Return: $D^{*}$
Algorithm 1 IcM Retrieval Process
6.3. TSG Retrieval
The get_tsg function is a critical component of our backend flow. Traditional retrieval-augmented generation (RAG) approaches typically use similarity-based searches, which compare user query embeddings with document chunk embeddings or employ term-frequency techniques (Robertson et al., 2009). However, empirical analyses show that these methods often underperform with highly technical documents, as embeddings do not capture the subtle distinctions within technical terms effectively. Although recent advancements in supervised retrieval models, such as RAFT (Zhang et al., 2024b), AR2 (Zhang et al., 2022), and ARG (Sachan et al., 2023), have improved retrieval accuracy, they necessitate the development and fine-tuning of deep neural networks (DNNs). These approaches are impractical for our self-hosting framework in DECO, as they require specialized machine learning expertise and additional infrastructure for training and hosting ML models, which may not be readily available to different teams.
To address these challenges, we developed a hybrid search algorithm, leveraging an advanced preprocessing pipeline detailed in Section 4, eliminating the need for additional ML training while enhancing both retrieval accuracy and computational efficiency. Building on the NL2SearchQuery pipeline, we introduce multiple search queries that run in parallel, we extract the following arguments for the search query:
  • Rephrased Query: User queries are rephrased by the planner to generate search arguments. Searches are conducted using both the original user query and the generated search arguments. For search query involving incidents, we generate additional “search hint” based on the incident summary, similar to HyDE (Zhao et al., 2023; Gao et al., 2022).
  • Search Field: Both “title” and “contents”.
  • Search Method: Multi-Field Hybrid Search based on reciprocal rank fusion (RRF) (Cormack et al., 2009; Microsoft Corporation, 2024) is used to combine the similarity score from all the search fields.
We also introduce several additional techniques to enhance retrieval results. A new Document Filtering step refines search results by analyzing the similarity scores of top-ranked documents during the re-ranking stage. If the highest-ranked document(s) show a significant margin over the following results, only these top documents are retained. This approach leverages the observation that a small set of highly relevant TSGs is often sufficient to resolve issues, reducing the inclusion of overly detailed or irrelevant documents and saving tokens for the final LLM calls (see Section 8.1).
Additionally, during preprocessing, teams provide concise Repository Descriptions for each document source. These descriptions are included with the retrieved documents to enhance the LLM’s contextual understanding, improving the quality of the final LLM call within the agent.
Future enhancements include incorporating advanced retrieval algorithms like HyDE and Reverse HyDE (Gao et al., 2022; Zhao et al., 2023), generating enriched matching fields, and exploring the integration of graph structures into the index (Anonymous, 2024a) and leveraging feedback from historical conversations to dynamically refine search rankings (Anonymous, 2024b).
6.4. Code Retrieval
A similar retrieval mechanism is also employed in the get_code skill for code search based on LA-RAG using the NL2SearchQuery algorithm. With few-shot examples, the arguments learned based on the NL2SearchQuery module include:
  • Search Text: Reformulated search elements drawn from the user’s question, maintaining the user’s intent and associated context, not merely the keywords.
  • Search Field: A list specifying the scope of the search, which can include one or more of the following values:
    1. (1)
      title”: Searches the codebase by title, useful for locating where specific elements like methods or classes are defined.
    2. (2)
      description”: Searches the codebase by the LLM generated description, applicable when user asks for explanation of the code.
    3. (3)
      content”: Searches the codebase by content, applicable when seeking definitions of specific methods or classes.
    4. (4)
      reference”: Searches the codebase by reference, targeting locations where specific methods or classes are referenced.
    5. (5)
      All: Searches across all vector indexes of the codebase.
7. Frontend
The frontend of DECO is built around the Azure Bot, which leverages the Azure Bot Framework V4 and is deployed on an Azure Web Application (see Figure 7). This setup coordinates API calls, manages messaging (interacting with the backend REST API), retains state in cloud storage, and supports user interface enhancements through frontend plugins such as Adaptive Cards (Microsoft, 2024a), allowing users to modify generated Kusto queries. Authentication is handled via Microsoft Entra ID (see Figure 7). The web application connects users to the backend Azure Bot through direct line web channels, which communicate with the Azure Bot API. This web channel seamlessly integrates with the IcM portal, pre-populating inputs like ticket IDs for backend processing. Additionally, the bot can be accessed via Microsoft Teams through a published manifest. The web application also transmits system telemetry to a reporting dashboard for monitoring. Detailed interactions between the components are shown in Figure 7.
Figure 7. Frontend components
7.1. Plugin
The frontend “Plugins” support various types of user interfaces required by certain skills. For example, for the kql.generator skill, the backend output includes plugin data that triggers a frontend plugin, which displays a special widget (an Adaptive Card) for the user. This widget allows users to select Kusto clusters from a dropdown menu, edit the query in the text box, use checkboxes to choose which queries to execute, and click the “submit” button to send the query back to the system. The user input data is stored in a designated field and sent to the backend, which then triggers the kql.executor skill to execute the query on behalf of the user.
7.2. Access Control
Access control for the bot is managed through Azure Active Directory (AAD) groups (Microsoft, 2024h). Although the web API endpoint is publicly accessible (necessary for communication with Teams, direct link channels, and potential future integrations), the bot itself handles the authentication process. The bot uses OAuth 2.0 (Microsoft, 2024i), with Azure Active Directory as the identity provider. This setup involves registering the bot with Azure, allowing it to function as an application that requests access to user data.
7.3. Telemetry
The copilot emits telemetry data to a cloud storage (i.g., Log Analytics Workspace), making this data accessible to copilot owners. Table 1 shows the types of data that are logged.
Table 1. Types of telemetry collected
Telemetry Type Details
Logins Records of logins and login attempts, including user email addresses. These logs are also used to compute Monthly Active Users (MAU).
Conversation statistics Logs of conversation volumes, latencies, and backend REST API call failures. This data is used to calculate retry percentages and answer success rates.
Feedback (stars) User feedback in the form of star ratings (1-5 stars).
Feedback (text) Users have the option to submit textual feedback, along with their names, allowing for follow-up communication if needed.
Conversation details
Full conversation text storage is optional and can be enabled at the discretion of the onboarded team.
Stored conversation details are anonymized and not explicitly linked to any user, ensuring privacy. A thorough security and privacy review was conducted to ensure compliance with the company’s data retention policy, which is based on well-defined data categories.
A dashboard has been developed to display high-level metrics for selected DECO bots, contingent on the onboarded team’s consent to share their telemetry data (see Figure 1).
8. Evaluation
The evaluation system in DECO is designed to support the continuous improvement of the framework by (1) helping to triage the reasoning behind low-quality responses and (2) evaluating the performance of individual components to guide further development and optimization. The evaluation process is divided into two categories: offline evaluation and online evaluation.
8.1. Offline Evaluation
To support ongoing development of the copilot and evaluate changes such as prompt modifications and updates to retrieval algorithms or skill selection mechanisms, we developed an evaluation framework that allows testing in hypothetical scenarios without disrupting the service. The offline evaluation focuses on assessing the quality of copilot responses after code changes.
We observed that LLM evaluation often differs significantly from human judgments, especially those made by domain experts, in terms of relevance and usefulness. Using a sample of user ratings, we explored the feasibility of employing an LLM to rate copilot responses on a scale from 1 to 5 stars, aiming at simulating human judgment. When comparing LLM-generated ratings with human ratings across 54 questions, we found a correlation of only 0.35 (see Figure 9), indicating substantial discrepancies between LLM and human evaluations. Manual reviews revealed that many responses deemed inaccurate by domain experts were perceived as plausible by non-experts due to subtle errors that only specialists could detect. This highlights the limitations of using LLMs as evaluators for assessing copilot responses, as discussed in (Liu et al., 2023). As a result, we prioritize using question-answer pairs from real user conversations with user ratings as benchmarks for evaluating copilot responses whenever feasible.
Evaluation Framework. We built an flexible offline evaluation framework (see Figure 8) for all the evaluation/test suites. The framework connects to the telemetry database using managed identity, allowing for the download of extensive real user queries, bot outputs, and user feedback. A curated set of highly user-rated question-answer pairs serves as the “golden answer” in the evaluation process. The framework features a runner that uses a configuration file to define which evaluators to include in the evaluation pipeline. Each evaluator, inheriting from a base class, implements three main functions:
  • Data Generator: Prepares input data from telemetry or predefined sources.
  • Execution Function: Runs the generated input data through specific components of the chat flow.
  • Evaluation Function: Produces evaluation metrics for each processed input.
  • Report Generator: Aggregates results from individual samples into a detailed dataset-level report, offering insights into chat flow performance.
Figure 8. Offline evaluation framework
Users can extend the framework by adding more evaluators using the provided templates. The evaluation pipeline is integrated into the Azure Pipeline and triggered by pull requests (PRs).
Evaluators. In this paper, we discuss four commonly used evaluators: the Planner Evaluator, the TSG Retrieval Evaluator, the Incident Retrieval Evaluator, and the Similarity Evaluator.
Figure 9. Comparison of human feedback with LLM-predicted ratings
Planner Evaluator
To evaluate the planner’s accuracy, we designed a series of questions intended to trigger different skills. For each question, a “golden answer” is provided, specifying the correct skills that should be invoked. The evaluation process for this module consists of the following steps:
  • Executing the planner module for each question.
  • Retrieving the list of skills invoked by the planner.
  • Comparing the invoked skills with the golden list. Precision and recall are calculated by comparing the set of selected skills with the golden set. Additionally, for each user question, we compute a Boolean metric named coverage, which indicates whether all required skills from the golden list are included in the triggered skills.
Whenever new skills/agents are added, we conduct experiments with this evaluator to ensure that the current planner can accurately select the skill when appropriate. If not, we fine-tune the skill’s description, the planner’s prompt, or the few-shot examples, to reach the optimal performance.
Figure 10 shows the evaluation results for different planner algorithms discussed in Section 5. Due to the inherent uncertainty in LLM calls, each setting includes five experimental runs, each with 18 representative user questions and verified required skill sets. We compare the use of few-shot vs. zero-shot examples and flat vs. agentic planners. We observe significant improvements when few-shot examples are used, particularly when incorporating a agentic structure for selecting the best skills with good consistency. In production, we also observed this behavior consistently.
Figure 10. Evaluation of different planner algorithms.
TSG Evaluator
The TSG Evaluator is designed to assess the accuracy of documentation chunks retrieved by the TSG retrieval process that are subsequently referenced in the copilot’s responses, i.e., by the get_tsg skill. The first “golden” set of TSG documentation is compiled from real user interactions, focusing only on responses rated four or five stars based on the “reference list” at the end of the copilot’s response, which includes clickable URLs to the cited documents as directed by the prompt. The second “golden” set is generated by the LLM, which creates synthetic questions for each documentation chunk, allowing us to identify the corresponding TSG documentation that answers those questions, similar to the self-supervised learning (Wang and Castro Fernandez, 2023).
The evaluation process involves the following steps:
  • Extracting the golden TSG list from conversations with highly rated responses or synthetic data, denoted as $G$.
  • Running the get_tsg module for each query to produce a new set of documents, referred to as $Q$. This step yields deterministic results without LLM calls. Optionally, the entire workflow can be executed to capture the final reference list generated by the LLM, denoted as $Q^{*}$.
  • Calculating precision and recall metrics by comparing $G$ and $Q^{*}$:
    (2) $\text{Precision}=\frac{|G\cap Q^{*}|}{|Q^{*}|},\quad\text{Recall}=\frac{|G\cap Q^{*}|}{|G|}.$
  • Assessing coverage by comparing $G$ with $Q$ to evaluate how well the retrieved documents ($Q$) “cover” the golden set ($G$):
    (3) $\text{Coverage}=\frac{|G\cap Q|}{|G|}.$
Figure 11 compares various retrieval algorithms as discussed in Section 6. The label “FT” indicates the use of document filtering, “RF” represents query rephrasing, and “HB” denotes hybrid semantic search. We tested 65 questions where the developer team has access to, with 28 sourced from real user data rated above four stars and the remaining 37 generated by LLM based on frequently retrieved document chunks. The results show that hybrid search (“HB-*”) significantly improves recall compared to other methods. Document filtering effectively reduces context length without compromising recall. Although query rephrasing does not notably enhance retrieval accuracy for these specific examples, it proves extremely useful for questions related to Incident Management tickets or those that are poorly formatted by users. For incident-related questions, where the raw query is combined with the incident ticket summary, search results are often suboptimal. However, rephrasing—similar to the HyDE process—enables the generation of optimized search text based on the ticket summary, improving document retrieval. Since generating ticket summaries involves an LLM call, the rephrasing process can be parallelized, avoiding additional latency on the critical path.
Figure 11. Comparison of different retrieval algorithms.
Figure 12 illustrates the comparison of context length and average recall across three most recent embedding models from Azure OpenAI (OpenAI, 2024): “text-embedding-3-large,” “text-embedding-3-small,” and “text-embedding-ada-002”, applied to the real user queries. The x-axis represents the configuration settings for the maximum number of documents retrieved. The results show that the “text-embedding-3-large” model, with similar context lengths, yields more accurate retrieval outcomes.
(a) Context length
(b) Average recall
Figure 12. TSG evaluator results.
We conclude that even within the current framework, configuration tuning, including the choice of search algorithm (e.g., hybrid vs. vector search), selection of search fields (e.g., title or content), and embedding models, significantly impacts retrieval performance.
Incident Retrieval Evaluator
To assess IcM similarity, we developed a detailed evaluation metric using LLMs by constructing a comprehensive prompt that compares various aspects of incident summaries (e.g., title, mitigation steps, and recorded root cause). For each pair of IcM summaries, we generate a similarity score of “Low”, “Medium”, or “High”. This allows us to evaluate the similarity between the summary of an incident in a user input that it and the summaries of retrieved incidents based on any search for similar incidents.
The evaluation process involves the following steps:
  • Collecting incident IDs referred to in real user queries.
  • Extracting the summary of each incident.
  • Running our get_icm module to extract corresponding historic similar incident summaries for each input.
  • Assessing the similarity between the input incident and the retrieved incidents.
We executed the evaluator against 100 sampled user questions. We compare these results against the baseline “Similar Incident API” developed by the IcM portal. As shown in Figure 13, our incident retrieval method significantly outperforms the baseline, retrieving more incidents with a “High” similarity score. Furthermore, our customized re-ranking algorithm (discussed in Section 6) further improves performance, achieving a “High” similarity score for 97 out of 100 incident pairs.
Figure 13. Evaluation of similar incident retrieval methods with and without the re-ranking algorithm compared with the default search algorithm for IcM portal.
Similarity Evaluator
To maintain consistency in response quality following code modifications, we rerun the copilot to generate new responses and benchmark them against historically high-rated answers. Specifically, we used LLM to assess the similarity between two answers to the same question, using a 1-5 star rating scale. Each newly generated response aims for a similarity rating of at least 4.0 when compared to prior highly-rated answers.
8.2. Online Evaluation
The online evaluation assesses the direct interactions between the copilot and users, producing performance indicators for actual conversations. In addition to user feedback ratings, we employ the following metrics: (1) relevance of the generated answer, scored from 1 to 3, (2) relevance of the retrieved documents, scored from 1 to 3, and (3) groundedness of the answer based on the retrieved documents, scored from 0 to 1. Each metric is evaluated using a specific prompt and an LLM to generate a numeric score. Additionally, an AML endpoint has been deployed to asynchronously collect supplementary telemetry data alongside the primary backend flow. We have categorized the results into five categories to clarify these scores for users (see Table 2).
Table 2. Answer categories
Category Details
Relevant-Grounded Answers with high relevance and grounded in documents.
Relevant-General Highly relevant answers based on model’s foundational training, not necessarily document-based.
Partially
Relevant-Grounded
Answers that are partially relevant and grounded in some documents.
Document Issue Responses with low relevance due to the absence of appropriate documents.
Grounding Issue Responses where relevant documents were found, but the answers were not well-grounded in the documented evidence.
For telemetry data collected through online evaluation (enabled by a subset of teams), we analyzed over 3,000 messages and categorized the responses as shown in Figure 14 from 2025-01-01 to 2025-02-28. The copilot averages 2.2 interaction rounds per session, with 6 rounds at the $95^{\text{th}}$ percentile and 11 rounds at the $99^{\text{th}}$ percentile, indicating that users often engage in multi-turn conversations. The most common issue identified was Documentation Issues (23%), while hallucination cases (i.e., Grounding Issues) accounted for 0.1%. Among conversations flagged as having documentation issues, the root causes can be attributed to three potential factors:
  1. (1)
    Missing documentation – The relevant document does not exist in the knowledge base.
  2. (2)
    Retrieval skills not being invoked – The planner fails to trigger the appropriate retrieval skills.
  3. (3)
    Irrelevant retrieved documents – The retrieved documents do not align with the user query.
Among responses identified as documentation issues, 68% of these responses state upfront that “no relevant information was found”, while 27% provide a reference list of documents for user verification. This suggests that the LLM effectively detects irrelevant retrieved documents and appropriately disregards them.
Figure 14. Online evaluation category distribution
Combining user feedback, among the responses where users gave a low rating ($\leq$ 1 star), 44% were identified with documentation issues. Among these, 93% explicitly indicated that no relevant information was found. Other low-rating responses where users provided text feedback highlight potential areas for copilot improvement, such as missing functionalities (e.g., inability to execute code) and off-topic responses. These issues may stem from inaccuracies in documentation retrieval or ambiguous user input.
9. Related Work
For automating error triage of IcM tickets, AutoTSG (Shetty et al., 2022) constructs executable scripts using TSG documentation, machine learning, and program synthesis to identify relevant troubleshooting steps. Prior works have also focused on automatic debugging and diagnosis of SQL query execution (Zhou et al., 2024), Spark jobs (Lu et al., 2019, 2017), and general root cause analysis for incidents (Roy et al., 2024; Zhang et al., 2024c; Ahmed et al., 2023; Zhang et al., 2024a; Chen et al., 2023; Demarne et al., 2020). AskBrain (Chen et al., 2020) applies language models to automatically categorize and assign incidents to appropriate teams, reducing mitigation time. However, these efforts primarily target specific categories of issues.
To improve retrieval-augmented generation (RAG) accuracy, TILDE (Zhuang and Zuccon, 2021b, a) and related works (Nogueira et al., 2019) propose re-ranking documents using fine-tuned likelihood or embedding models, as well as pretrained relevancy models (Nogueira et al., 2020). ReSP (Jiang et al., 2024) iteratively retrieves documents using a “Reasoner” to assess the current retrieved documents, although this introduces significant latency. (Ma et al., 2024) fine-tunes the LLaMA model as both a dense retriever and a re-ranker in a multi-stage text retrieval pipeline, achieving improved retrieval performance. Retrieval-augmented methods have also been applied to code repositories for code completion tasks, such as RepoCoder (Zhang et al., 2023) and GraphCoder (Liu et al., 2024). However, these systems typically focus on generating code completions from partially written code, rather than handling natural language queries.
Recent research has also explored agentic frameworks to manage complex workflows involving multiple tools and skills. For example, systems such as AutoGen (Wu et al., 2023) and Letta (Packer et al., 2023) introduce multi-agent coordination patterns for task planning and execution, primarily in research or prototype environments. Gorilla (Patil et al., 2023) focuses on large-scale API selection by prompting language models to generate function calls from extensive catalogs, demonstrating the potential of LLMs for dynamic skill invocation. LlamaIndex (Liu and contributors, 2022) provides a general-purpose framework for connecting LLMs with external knowledge sources. In contrast, DECO adopts a lightweight agentic framework optimized for production-scale deployments, decoupling backend and frontend interactions. This design improves scalability and adaptability, particularly as the number of available skills grows in enterprise environments.
10. Conclusion
In this paper, we introduced DECO, a framework for end-to-end management and deployment of copilots tailored to software engineering tasks. DECO integrates preprocessing pipelines with tightly connected front-end and back-end modules to support multi-turn interactions with users. Leveraging historical incident tickets, DECO bridges knowledge gaps by enriching documentation with insights from past engineering practices. Its hierarchical skill selection with agentic design dynamically selects the most relevant context and tools, using advanced retrieval algorithms to surface pertinent incidents, documents, and code snippets. A continuous evaluation pipeline monitors retrieval performance and response quality. Deployed across Microsoft, DECO now serves over 500 Monthly Active Users (MAUs), reducing time spent on incident management and improving engineering workflows. Future work will focus on refining retrieval algorithms, improving document relevance through user feedback, and enhancing memory management to handle longer chat histories more efficiently.